OpenStudy (anonymous):

hi i need help with 2(x+1)^2-3(x+1)

OpenStudy (anonymous):

what do you need to do with it? multiply out maybe?

OpenStudy (anonymous):

\[2(x+1)^2-3(x+1)=2(x+1)(x+1)-3x-3\] is a start then use the distributive property again for the first part

OpenStudy (anonymous):

yes its part of a function problem f(x+1)= 2x^2-3x

OpenStudy (anonymous):

\[2(x+1)(x1)-3x-3=2(x^2+2x+1)-3x-3=2x^2+4x+2-3x-3\] so finally you get \[2x^2+x-1\]

OpenStudy (anonymous):

where is the (x1) from did u mean (x+1)?

OpenStudy (anonymous):

yes a typo

OpenStudy (anonymous):

ok thank u